OpenStudy (anonymous):

What is the slope of the line passing through (-3, 5) and (5, -3)? slope = -1 slope = 0 slope = 1 no slope

OpenStudy (anonymous):

slope = (y2-y1) / (x2-x1) = (-3-5) / (5+3) = -8/8 = -1

OpenStudy (amorfide):

 substitute your values